OpenStudy (anonymous):

In order for a satellite to remain in a circular orbit above the same spot on Earth, the satellite must be 35,800 km above the equator. Write an equation for the orbit of the satellite. Use the center of Earth as the origin and 6400 km for the radius of the Earth.

OpenStudy (anonymous):

Satellite is 35,800 km above the equator and 35,800 + 6,400 = 42,200 km from centre of the earth Orbit of satellite will be along the circular path: x² + y² = 42,200²

OpenStudy (anonymous):

or Taking (0,0) as the center of the earth, and of the orbit. General cartesian eqn for the circle : (x-a)^2 + (y-b^)2 = r^2 But (a,b) is (0,0), as stated above So : x^2 + y^2 = r^2 What is r ? r = 6400 + 35800 = 42200 So the eqn is : x^2 + y^2 = 42200 ^2

OpenStudy (anonymous):

do you get it or no